OpenStudy (anonymous):

If f(x) = |x| + 9 and g(x) = –6, which describes the value of (f + g)(x)? (f + g)(x) greater than or equal to 3 for all values of x (f + g)(x) less than or equal to 3 for all values of x (f + g)(x) greater than or equal to 6 for all values of x (f + g)(x) lless than or equal to 6 for all values of x

OpenStudy (anonymous):

(f + g)(x) = f(x) + g(x) = |x| + 9 + (-6) = |x| + 3 |x| is either equal to zero or greater than zero, so (f + g)(x) greater than or equal to 3 for all values of x

OpenStudy (anonymous):

Gracias! Had that in mind. Wasn't so sure though.